• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:04:45 on December 13, 2006

    Entry number 190664

    keyword=Shift Summary

    Shift Summary Day 12/13/06
    John LeRose (SL)
    Bob Michaels (TO)

    Program is to continue taking data at KIN #7
    Left Arm: 2.849 GeV/c, 28.792°
    Right Arm: P=0.7843 GeV/c, 60.807°

    09:55 Observe that the right NMR is not updating. Magnet current
    is stable and updating.
    Since, we will lose the beam at 10:00 for beam studies will
    reboot the IOC then.
    10:01 Beam off for studies
    10:13 ARM arrives for controlled access survey and Informs us
    that the beam was off for beam studies. Paged Ed to let him know.
    10:39 Reboot of right dipole NMR IOC is successful
    13:00 Tech's and Bob and Alex are out of the Hall. Rebooted
    left arm controls in order to get left dipole to set after
    tech's worked on it.
    13:22 Mike Seely goes into the Hall to check some target controls
    13:40 Mike and Co. are out of the Hall
    14:01 MCC says no beam till after 16:00. Jack goes into the Hall to
    work on the right arm reboot switch.
    15:05 Jack is finished in the Hall

    run start/stop comments
    1478 07:25/10:13 Beam stopped at 10:01 for beam studies (nobody told us)